• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Burtin

    Log entry time 23:55:49 on October15,2005

    Entry number 156996

    keyword=evening shift summary

    16:30 access to reset right arm magnet controls
    16:45 rebooted the magnet control ioc of the right arm
    16:46 right arm septum tripped
    17:00 Brian fixed the sync problem in the right arm
    17:30 both dipoles have tripped
    19:45 no hope to fix right dipole today - take data with one arm. end of access.
    20:45 injector problem no beam yet
    21:15 beam back 5 muA
    21:20 beam at 20 muA
    21:25 beam at 40 muA
    21:35 requested 60 muA
    21:50 MCC called because they cannot do better than 40 muA because of trip in the ion chamber in the left dump
    22:10 after David's intervention they raise the trip limit and got us to 60 muA
    22:30 adjust detector HV to accouont for these conditions.

    5655 - production run with 1 arm 60 muA.

    23:00 we will turn the right septum on. ask for beam off and wait that the temperature of the coil goes below the limit of 7.53K.

    23:10 start ramping up the current in the R septum to 316.92 and set the L septum at 323.61

    23:45 beam back

    5656 - production run with 1 arm 58 muA